# Improving Stance Detection by Leveraging Measurement Knowledge from Social Sciences: A Case Study of Dutch Political Tweets and Traditional Gender Role Division

### v2

@@ Abstract
Stance detection (SD) concerns automatically determining the viewpoint (i.e., in favour of, against, or neutral) of a text’s author towards a target. SD has been applied to many research topics, among which the detection of stances behind political tweets is an important one. In this paper, we apply SD to a dataset of tweets from official party accounts in the Netherlands between 2017 and 2021, with a focus on stances towards traditional gender role division, a dividing issue between (some) Dutch political parties. To implement and improve SD of traditional gender role division, we propose to leverage an established survey instrument from social sciences, which has been validated for the purpose of measuring attitudes towards traditional gender role division. Based on our experiments, we show that using such a validated survey instrument helps to improve SD performance.
@@ 1 Introduction
Stance detection (SD) is an important natural language processing (NLP) task. SD aims to automatically determine the viewpoint of a text’s author towards a target (Mohammad et al., 2016). This viewpoint, also known as stance, can be in favour of, against or neutral. Typical targets include various social and political topics (e.g., abortion, feminism), political figures (e.g., Trump, Clinton) and events (e.g., referendums) (Lahoti et al., 2018; Lai et al., 2020; Aldayel and Magdy, 2021). SD has been extensively applied to various research areas, including political discourse analysis (Gunhal et al., 2022; Lai et al., 2020). Political tweets, in particular, serve as a rich source of information for understanding political stances and attitudes, making SD particularly relevant in this domain.
@@
Detecting stances behind political tweets is a multifaceted task that involves analysing the linguistic nuances and contextual factors embedded within tweets to determine the viewpoint expressed towards a specific issue or target. Traditionally, this process has been approached using supervised learning methods, where a classifier is trained on a dataset of labelled tweets, where each tweet is associated with a predefined stance category (Li et al., 2021; Samih and Darwish, 2021; Lai et al., 2020). Unsupervised learning methods for SD also exist, which do not rely on labelled data, but instead, seek to identify latent patterns and structures within a dataset of unlabelled tweets (Gambini et al., 2023; Darwish et al., 2020; Stefanov et al., 2020). They typically employ techniques such as topic modelling, sentiment analysis, and distributional semantics to extract latent features from the tweets. These features are then analysed to identify clusters or groupings of tweets that share similar linguistic characteristics and semantic associations. These clusters are then assigned stance labels based on their perceived sentiment or topic orientation.
@@
Recently, SD is increasingly conceptualised as textual entailment recognition (TER) (Aldayel and Magdy, 2021; Mohammad et al., 2016). TER, a more generic natural language understanding task, concerns determining whether a hypothesis follows from a premise, where three outcomes are possible: “entailment”, “contradiction” and “neutral” (Dagan and Glickman, 2004). By reformulating SD as judging whether a text (i.e., a premise) entails a stance towards a target (i.e., a hypothesis), a SD problem becomes a TER one. This reformulation offers two potential benefits. First, TER-based models are not target-specific and may be applied to unseen targets. Second, TER datasets can be used for training SD models. Currently, labelled SD datasets are limited to a few pre-defined targets and languages (Aldayel and Magdy, 2021). In contrast, TER datasets are not target-specific and are available in more languages. Therefore, we can use TER datasets (in addition to SD ones) for model training when considering SD as TER. Research has shown that training on additional TER datasets improves downstream SD performance (Schiller et al., 2021).
@@
In this paper, we investigate the application of SD as TER to a dataset of tweets from political parties and politicians in the Netherlands between 2017 and 2021, focusing specifically on stances towards traditional gender role division. This issue holds significant political and social implications, often dividing various Dutch political parties. To further improve SD performance and capture nuanced stances, we propose utilising a validated survey instrument from social sciences specifically designed to measure attitudes towards traditional gender role division. Our experiments demonstrate that incorporating this instrument indeed leads to improved SD performance. To the best of our knowledge, our study is the first to incorporate survey instruments for SD.
@@ 2.1 Stance Detection as Textual Entailment Recognition
TER, as mentioned in section 1, concerns judging whether a hypothesis ($h$) can be inferred from a premise ($p$). If yes, we conclude that $p$ entails $h$ ($p\Rightarrow h$). Typically, three outcomes are possible, including “entailment”, “contradiction” and “neutral”. Note that the relation between $p$ and $h$ is directional, where $p\Rightarrow h$ does not imply $h\Rightarrow p$.
The connection between the two tasks, TER and SD, becomes clear when we reformulate SD as judging whether a text ($t$) entails a stance to a target (i.e., $t\Rightarrow stt$). For instance, say the text $t$ is “I appreciate more women in government positions”, the target is “traditional gender role division”, and the stance is “in favour of”. Then, we can formulate the stance to the target $sst$ as “I am in favour of traditional gender role division”. As such, the task becomes judging whether “I appreciate more women in government positions” entails, contradicts or is neutral towards “I am in favour of traditional gender role division”. The three possible TER outcomes “entailment”, “contradiction” and “neutral” correspond to the three SD outcomes: “in favour of”, “against” and “neutral”.
One benefit of considering SD as TER is that a model can be easily trained on data involving multiple targets and can incorporate information about the targets during training. This may allow the model to predict stances towards unseen targets, also known as zero-shot learning (Yin et al., 2019). Another benefit is that TER datasets can be used for model training. For languages where existing SD datasets are available, additional TER datasets may help to further improve model performance. When SD datasets do not exist for a language but TER datasets do, we can still train a model based on TER datasets to predict stances. It is, however, important to note that current TER datasets are quite different from SD datasets (Bowman et al., 2015; Williams et al., 2018). The former mainly consists of descriptions of concrete scenes (e.g., “two dogs are fighting”) which assume little world knowledge (e.g., social and political topics), while the latter are typically user-generated texts (e.g., social media posts) and concern diverse targets that can be abstract and ambiguous. This data discrepancy may limit the benefit of modelling SD as TER.
@@ 2.2 Stance Detection as Attitudinal Measurements in Social Sciences
In social sciences, stance detection is treated as a measurement problem and typically relies on the use of survey instruments. For instance, to measure stances towards traditional gender role division, one can ask a participant to indicate on a Likert scale from 1 to 5 how much they agree with the following statement: “Overall, family life suffers the consequences if the mother has a full-time job.” (Stam et al., 2014). The more they agree with this statement, the more they are in favour of the target (i.e., traditional gender role division).
Survey instruments need to be both “valid” and “reliable”. “Valid” means that they measure what they are supposed to, while “reliable” means that the measurements do not suffer from large random variations (Trochim et al., 2015). One common way to establish the validity of survey instruments is to ensure that they cover all the aspects of a given target. This is also known as “content validity” (Trochim et al., 2015). For instance, when measuring stances to traditional gender role division, it is important to cover relevant subdomains of this target like in career, childcare, housework and education (Walter, 2018). Furthermore, higher reliability can be achieved by averaging the responses to multiple survey instruments concerning the same target.
Established survey instruments exist for many SD targets (e.g., feminism (Frieze and McHugh, 1998), abortion (Hendriks et al., 2012), climate change (Dijkstra and Goedhart, 2012) and politicians (Martin et al., 2005)). They can potentially benefit the modelling of SD as TER. For instance, when using a model trained on TER datasets to detect stances towards traditional gender role division, the hypothesis can be as simple as “I am in favour of traditional gender role division”. This hypothesis, however, does not explicitly cover all the relevant subdomains of the target (i.e., in career, childcare, housework and education). Consequently, the model may fail to identify texts that entail these subdomain stances. We may mitigate this issue by using established survey items (that have demonstrated content validity) to construct multiple hypotheses. Averaging the model’s predictions across these multiple hypotheses can also lead to more reliable, less random results. Therefore, constructing multiple hypotheses based on established survey instruments for a given target may improve prediction validity and reliability.
@@ 3 Experiments
We investigate the approach of modelling SD as TER by applying it to a dataset of tweets from Dutch political parties and politicians in the Netherlands, with the goal to measure stances towards traditional gender role division.
@@ 3.1 Datasets
We use Twitter’s API to collect a dataset of tweets from the official Twitter accounts of 13 major Dutch political parties (see Table 1) and the accounts of 10 politicians per party. We collect up to 3,200 most recent tweets per account (API limit) and in total, 311,362 tweets. We lower-case the tweets, remove urls and special characters, leave out tweets shorter than five words, and retain only the ones posted between 2017 and 2021. The resulting dataset contains 247,902 tweets. To reduce the number of tweets unrelated to the target of interest, we further apply a simple filtering strategy using keywords related to the target, [^1] The Dutch equivalent of “woman”, “man”, “mother”, “father”, “boy” and “girl”: “vrouw”, “man”, “moeder”, “vader”, “jongen”, “meisje”. leading to a much smaller final dataset of 2,601 tweets.
Previous research suggests that a party’s stance towards an issue aligns with the average opinion of its voters (Ibenskas and Polk, 2022). This inspires us to use the 2017-2021 data from the LISS panel [^2] https://www.lissdata.nl/about-panel to provide ground truth labels on the party level. The LISS panel is a representative sample of Dutch households that participate in monthly surveys about various topics like health, work and political views. The LISS data contains responses to 11 established survey items about stances towards traditional gender role division (see Section B), as well as the party for which the participant last voted. We group the participants by the party they voted for and average their responses according to each participant’s respective sampling weight across the 11 survey items for a given year and across all years. These scores are used to rank the parties in terms of their year(s)-specific stances towards traditional gender role division (see Table 1). For tweet-level ground truth labels, we rely on manual annotation (see Section 3.2 for more information, and Appendix C for annotated examples).
@@
Furthermore, we use the SICK-NL dataset (Wijnholds and Moortgat, 2021), which is the Dutch translation of the original English SICK dataset (Marelli et al., 2014). It contains 9,840 pairs of sentences that describe concrete scenes (e.g., “Two dogs are fighting”), including 5,595 neutral pairs, 1,424 contradiction pairs and 2,821 entailment pairs. To the best of our knowledge, SICK-NL is the only Dutch TER dataset. Also, no labelled SD dataset exists for the Dutch language.
3.2 Experiment 1: Zero-shot SD with BERTje
Due to the lack of labeled Dutch SD datasets, we choose to fine-tune a TER classifier on the available SICK-NL dataset, and subsequently apply the trained classifier to perform zero-shot prediction of stances towards traditional gender role division. Specifically, we fine-tune BERTje, a pretrained Dutch BERT model (de Vries et al., 2019), on the SICK-NL dataset, using 9,345 pairs for training and the rest for validation. We use the same hyperparameters as Wijnholds and Moortgat (2021). See Appendix A for more computational details. Next, we apply the fine-tuned BERTje model to our dataset of Dutch political tweets, where the goal is to predict whether a tweet entails a hypothesis (i.e., a stance towards traditional gender role division). We use two types of hypotheses. The first is a simple statement (in Dutch): “I am in favour of traditional gender role division”. The second consists of the 11 survey items about traditional gender role division from the LISS panel (also in Dutch). See Section B for more information about the construction of hypotheses.
We evaluate the SD performance at both the individual tweet level and the party level. Due to the large number of tweets in the dataset and lack of resources for manual annotation of all the tweets, we focus on evaluating only the top $k$ tweets with the highest predicted entailment scores, where $k={10,50,100}$. The more out of the $k$ tweets are, indeed, true entailments, the more evidence that the model has encoded information useful for (zero-shot) stance prediction.
@@ 3.2 Experiment 1: Zero-shot SD with BERTje
Results
Table 2 summarises the results on top-$k$ precision and Spearman’s $\rho$. For the rows pertaining to top-$k$ precision, the scores are the percentages of tweets correctly classified as entailments. We can make two observations. First, on the tweet level, the best precision scores consistently occur when survey items are used, indicating that the use of survey items to construct multiple hypotheses can be beneficial for SD. Second, however, the Spearman’s $\rho$ scores paint a different picture: all the scores are very far from 1, indicating poor performance on the party level. These findings suggest that our Bertje model encodes useful information for zero-shot prediction of an unseen stance target, but only when survey items are used.
@@ Results
Spearman’s $\rho_{10}$0.04 -0.07
Spearman’s $\rho_{50}$ 0.06 -0.07
Spearman’s $\rho_{100}$ 0.15 -0.08
Spearman’s $\rho_{all}$ 0.13 -0.09
3.3 Experiment 2: Zero-shot SD with GPT-3.5 Turbo
To further validate our findings from Experiment 1, we conduct a second experiment where we use prompts with GPT-3.5 Turbo for zero-shot SD of traditional gender role division. We use a balanced dataset of 200 political tweets that were labelled during the annotation procedure of Experiment 1. We use the OpenAI API with GPT-3.5 Turbo, with the following prompt template, where {premise} and {hypothesis} refer to two separate string variables:
@@
We use an English template because in this case, GPT-3.5 Turbo gives responses that are consistent with the requirement of the prompt (i.e., returning only “entails”, “contradicts” or “neutral”). In contrast, using a Dutch template would throw off the model, leading it to return sentences (e.g., “The premise entails the hypothesis”; “There is no relationship between the premise and the hypothesis”). To improve reproducibility, we set the temperature parameter to be 0 and the random seed to 1.
@@
Using only a simple hypothesis, we obtain an accuracy score of 0.58; using all the survey items, however, we obtain a higher accuracy score of 0.725. This further suggests evidence for the efficacy of using survey items for SD.

### v3

@@ Abstract
Stance detection concerns automatically determining the viewpoint (i.e., in favour of, against, or neutral) of a text’s author towards a target. Stance detection has been applied to many research topics, among which the detection of stances behind political tweets is an important one. In this paper, we apply stance detection to a dataset of tweets from official party accounts in the Netherlands between 2017 and 2021, with a focus on stances towards traditional gender role division, a dividing issue between (some) Dutch political parties. To implement and improve stance detection of traditional gender role division, we propose to leverage an established survey instrument from social sciences, which has been validated for the purpose of measuring attitudes towards traditional gender role division. Based on our experiments, we show that using such a validated survey instrument helps to improve stance detection performance.
@@ 1 Introduction
Stance detection is an important natural language processing (NLP) task. Stance detection aims to automatically determine the viewpoint of a text’s author towards a target [17]. This viewpoint, also known as stance, can be in favour of, against or neutral. Typical targets include various social and political topics (e.g., abortion, feminism), political figures (e.g., Trump, Clinton) and events (e.g., referendums) [12, 13, 1]. Stance detection has been extensively applied to various research areas, including political discourse analysis [9, 13]. Political tweets, in particular, serve as a rich source of information for understanding political stances and attitudes, making stance detection particularly relevant in this domain.
@@
Detecting stances behind political tweets is a multifaceted task that involves analysing the linguistic nuances and contextual factors embedded within tweets to determine the viewpoint expressed towards a specific issue or target. Traditionally, this process has been approached using supervised learning methods, where a classifier is trained on a dataset of labelled tweets, where each tweet is associated with a predefined stance category [14, 18, 13]. Unsupervised learning methods for stance detection also exist, which do not rely on labelled data, but instead, seek to identify latent patterns and structures within a dataset of unlabelled tweets [8, 4, 21]. They typically employ techniques such as topic modelling, sentiment analysis, and distributional semantics to extract latent features from the tweets. These features are then analysed to identify clusters or groupings of tweets that share similar linguistic characteristics and semantic associations. These clusters are then assigned stance labels based on their perceived sentiment or topic orientation.
@@
Recently, stance detection is increasingly conceptualised as textual entailment recognition (TER) [1, 17]. TER, a more generic natural language understanding task, concerns determining whether a hypothesis follows from a premise, where three outcomes are possible: “entailment”, “contradiction” and “neutral” [3]. By reformulating stance detection as judging whether a text (i.e., a premise) entails a stance towards a target (i.e., a hypothesis), a stance detection problem becomes a TER one. This reformulation offers two potential benefits. First, TER-based models are not target-specific and may be applied to unseen targets. Second, TER datasets can be used for training stance detection models. Currently, labelled stance detection datasets are limited to a few pre-defined targets and languages [1]. In contrast, TER datasets are not target-specific and are available in more languages. Therefore, we can use TER datasets (in addition to stance detection ones) for model training when considering stance detection as TER. Research has shown that training on additional TER datasets improves downstream stance detection performance [19].
@@
In this paper, we investigate the application of stance detection as TER to a dataset of tweets from political parties and politicians in the Netherlands between 2017 and 2021, focusing specifically on stances towards traditional gender role division. This issue holds significant political and social implications, often dividing various Dutch political parties. To further improve stance detection performance and capture nuanced stances, we propose utilising a validated survey instrument from social sciences specifically designed to measure attitudes towards traditional gender role division. Our experiments demonstrate that incorporating this instrument indeed leads to improved stance detection performance. To the best of our knowledge, our study is the first to incorporate survey instruments for stance detection.
@@ 2.1 Stance Detection as Textual Entailment Recognition
TER, as mentioned in Section 1, concerns judging whether a hypothesis ($h$) can be inferred from a premise ($p$). If yes, we conclude that $p$ entails $h$ ($p\Rightarrow h$). Typically, three outcomes are possible: “entailment”, “contradiction” and “neutral”. Note that the relation between $p$ and $h$ is directional, where $p\Rightarrow h$ does not imply $h\Rightarrow p$.
The connection between the two tasks, TER and stance detection, becomes clear when we reformulate stance detection as judging whether a text ($t$) entails a stance to a target (i.e., $t\Rightarrow stt$). For instance, say the text $t$ is “I appreciate more women in government positions”, the target is “traditional gender role division”, and the stance is “in favour of”. Then, we can formulate the stance to the target sstsst as “I am in favour of traditional gender role division”. As such, the task becomes judging whether “I appreciate more women in government positions” entails, contradicts or is neutral towards “I am in favour of traditional gender role division”. The three possible TER outcomes “entailment”, “contradiction” and “neutral” correspond to the three stance detection outcomes: “in favour of”, “against” and “neutral”.
One benefit of considering stance detection as TER is that a model can be easily trained on data involving multiple targets and can incorporate information about the targets during training. This may allow the model to predict stances towards unseen targets, also known as zero-shot learning [26]. Another benefit is that TER datasets can be used for model training. For languages where existing stance detection datasets are available, additional TER datasets may help to further improve model performance. When stance detection datasets do not exist for a language but TER datasets do, we can still train a model based on TER datasets to predict stances. It is, however, important to note that current TER datasets are quite different from stance detection datasets [2, 25]. The former mainly consists of descriptions of concrete scenes (e.g., “two dogs are fighting”) which assume little world knowledge (e.g., social and political topics), while the latter are typically user-generated texts (e.g., social media posts) and concern diverse targets that can be abstract and ambiguous. This data discrepancy may limit the benefit of modelling stance detection as TER.
@@ 2.2 Stance Detection as Attitudinal Measurements in Social Sciences
In social sciences, stance detection is treated as a measurement problem and typically relies on the use of survey instruments. For instance, to measure stances towards traditional gender role division, one can ask a participant to indicate on a Likert scale from 1 to 5 how much they agree with the following statement: “Overall, family life suffers the consequences if the mother has a full-time job.” [20]. The more they agree with this statement, the more they are in favour of the target (i.e., traditional gender role division).
Survey instruments need to be both “valid” and “reliable”. “Valid” means that they measure what they are supposed to, while “reliable” means that the measurements do not suffer from large random variations [22]. One common way to establish the validity of survey instruments is to ensure that they cover all the aspects of a given target. This is also known as “content validity” [22]. For instance, when measuring stances to traditional gender role division, it is important to cover relevant subdomains of this target like in career, childcare, housework and education [23]. Furthermore, higher reliability can be achieved by averaging the responses to multiple survey instruments concerning the same target.
Established survey instruments exist for many stance detection targets (e.g., feminism [7], abortion [10], climate change [6] and politicians [16]). They can potentially benefit the modelling of stance detection as TER. For instance, when using a model trained on TER datasets to detect stances towards traditional gender role division, the hypothesis can be as simple as “I am in favour of traditional gender role division”. This hypothesis, however, does not explicitly cover all the relevant subdomains of the target (i.e., in career, childcare, housework and education). Consequently, the model may fail to identify texts that entail these subdomain stances. We may mitigate this issue by using established survey items (that have demonstrated content validity) to construct multiple hypotheses. Averaging the model’s predictions across these multiple hypotheses can also lead to more reliable, less random results. Therefore, constructing multiple hypotheses based on established survey instruments for a given target may improve prediction validity and reliability.
@@ 3 Experiments
We investigate the approach of modelling stance detection as TER by applying it to a dataset of tweets from Dutch political parties and politicians in the Netherlands, with the goal to measure stances towards traditional gender role division. Specifically, we conduct two experiments in a zero-shot stance detection setting. In the first experiment (Section 3.2), we fine-tune Bertje [5], a Dutch BERT model, using Dutch TER data. Furthermore, we conduct a second experiment (Section 3.3) to validate the results from the first experiment. Given the recent developments and promising applications of large generative language models in various NLP tasks, we prompt GPT-3.5 Turbo, a large language model, to perform zero-shot stance detection.
@@ 3.1 Datasets
We use Twitter (now X)’s API to collect a dataset of tweets from the official Twitter accounts of 13 major Dutch political parties (see Table 1) and the accounts of 10 politicians per party. We collect up to 3,200 most recent tweets per account (API limit) and in total, 311,362 tweets. We lower-case the tweets, remove urls and special characters, leave out tweets shorter than five words, and retain only the ones posted between 2017 and 2021. The resulting dataset contains 247,902 tweets. To reduce the number of tweets unrelated to the target of interest, we further apply a simple filtering strategy using keywords related to the target, [^1] The Dutch equivalent of “woman”, “man”, “mother”, “father”, “boy” and “girl”: “vrouw”, “man”, “moeder”, “vader”, “jongen”, “meisje”. leading to a much smaller final dataset of 2,601 tweets.
Previous research suggests that a party’s stance towards an issue aligns with the average opinion of its voters [11]. This inspires us to use the 2017-2021 data from the LISS panel [^2] https://www.lissdata.nl/about-panel to provide ground truth labels on the party level. The LISS panel is a representative sample of Dutch households that participate in monthly surveys about various topics like health, work and political views. The LISS data contains responses to 11 established survey items about stances towards traditional gender role division (see Section 0.B), as well as the party for which the participant last voted. We group the participants by the party they voted for and average their responses according to each participant’s respective sampling weight across the 11 survey items for a given year and across all years. These scores are used to rank the parties in terms of their year(s)-specific stances towards traditional gender role division (see Table 1). For tweet-level ground truth labels, we rely on manual annotation (see Section 3.2 for more information, and Appendix 0.C for annotated examples).
@@
Furthermore, we use the SICK-NL dataset [24], which is the Dutch translation of the original English SICK dataset [15]. It contains 9,840 pairs of sentences that describe concrete scenes (e.g., “Two dogs are fighting”), including 5,595 neutral pairs, 1,424 contradiction pairs and 2,821 entailment pairs. To the best of our knowledge, SICK-NL is the only Dutch TER dataset. Also, no labelled stance detection dataset exists for the Dutch language.
3.2 Experiment 1: Zero-shot Stance Detection with BERTje
Due to the lack of labeled Dutch stance detection datasets, we choose to fine-tune a TER classifier on the available SICK-NL dataset, and subsequently apply the trained classifier to perform zero-shot prediction of stances towards traditional gender role division. Specifically, we fine-tune BERTje, a pretrained Dutch BERT model [5], on the SICK-NL dataset, using 9,345 pairs for training and the rest for validation. We use the same hyperparameters as [24]. See Appendix 0.A for more computational details. Next, we apply the fine-tuned BERTje model to our dataset of Dutch political tweets, where the goal is to predict whether a tweet entails a hypothesis (i.e., a stance towards traditional gender role division). We use two types of hypotheses. The first is a simple statement (in Dutch): “I am in favour of traditional gender role division”. The second consists of the 11 survey items about traditional gender role division from the LISS panel (also in Dutch). See Section 0.B for more information about the construction of hypotheses.
We evaluate the stance detection performance at both the individual tweet level and the party level. Due to the large number of tweets in the dataset and lack of resources for manual annotation of all the tweets, we focus on evaluating only the top $k$ tweets with the highest predicted entailment scores, where k=10,50,100k={10,50,100}. The more out of the $k$ tweets are, indeed, true entailments, the more evidence that the model has encoded information useful for (zero-shot) stance prediction.
@@ 3.2 Experiment 1: Zero-shot Stance Detection with BERTje
Spearman’s $\rho_{10}$ - -0.04 -0.07
Spearman’s $\rho_{50}$ - 0.06 -0.07
Spearman’s $\rho_{100}$ - 0.15 -0.08
Spearman’s $\rho_{all}$ - 0.13 -0.09
Results
Table 2 summarises the results on top-$k$ precision and Spearman’s $\rho$. For the rows pertaining to top-$k$ precision, the scores are the percentages of tweets correctly classified as entailments. We can make two observations. First, on the tweet level, the best precision scores consistently occur when survey items are used, indicating that the use of survey items to construct multiple hypotheses can be beneficial for stance detection. Second, however, the Spearman’s $\rho$ scores paint a different picture: all the scores are very far from 1, indicating poor performance on the party level. These findings suggest that our Bertje model encodes useful information for zero-shot prediction of an unseen stance target, but only when survey items are used.
3.3 Experiment 2: Zero-shot Stance Detection with GPT-3.5 Turbo
To further validate our findings from Experiment 1, we conduct a second experiment where we use GPT-3.5 Turbo, a large language model, to perform zero-shot stance detection of traditional gender role division. We use a balanced dataset of 200 political tweets that were labelled during the annotation procedure of Experiment 1. We use the OpenAI API to prompt GPT-3.5 Turbo, with the following prompt template, where {premise} and {hypothesis} refer to two separate string variables:
@@ Results
We use an English template because in this case, GPT-3.5 Turbo gives responses that are consistent with the requirement of the prompt (i.e., returning only “entails”, “contradicts” or “neutral”). In contrast, using a Dutch template would throw off the model, leading it to return sentences (e.g., “The premise entails the hypothesis”; “There is no relationship between the premise and the hypothesis”) and sometimes, responses in other languages like English. To improve reproducibility, we set the temperature parameter to be 0 and the random seed to 1.
@@
Using only a simple hypothesis, we obtain an accuracy score of 0.580; using all the survey items, however, we obtain a higher accuracy score of 0.725. This further suggests evidence for the efficacy of using survey items for stance detection.